A tomato gun can fire a tomato at 17.2 km/h and is aimed at an very innocent bystander. Upon impact, if the tomato stops within 6.6cm, find the acceleration experienced by the tomato.
1
Expert's answer
2019-10-22T10:20:48-0400
"v^2=2as"

"\left(\frac{17.2}{3.6}\right)^2=2(0.066)a"

"a=173\frac{m}{s^2}"